2. We take a random sample of 250 Maryland police officers and find that their average annual salary is $42,500, with a standard deviation of $6,000. Construct a 98% confidence interval around your point estimate. Interpret your results.


1
Expert's answer
2021-12-12T17:42:37-0500

The critical value for "\alpha = 0.02" and "df = n-1 =249" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} = 2.341417"

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(42500-2.341417\times\dfrac{6000}{\sqrt{250}},"

"42500+2.341417\times\dfrac{6000}{\sqrt{250}})"

"=(41611.495, 43388.505)"

Therefore, based on the data provided, the 98% confidence interval for the population mean is "41611.495 < \mu < 43388.505," which indicates that we are 98% confident that the true population mean "\mu" is contained by the interval "(41611.495, 43388.505)."